OpenStudy (anonymous):

Use complete sentences to analyze the graph of the function f(x) = log x base 1/5. Part 1 [4 points]: Describe the domain, range, and general shape of this graph. Part 2 [4 points]: Using complete sentences, explain how plotting specific points helps graph the function and note any critical points such as its intercepts.

OpenStudy (dumbcow):

The log function has vertical asymptote at x=0 x-intercept at x=1 Because its base 1/5, when x =1/5, y = 1 that should give the general shape of the graph

OpenStudy (anonymous):

also you might note that since the base is less than 1, this curve decreases